OpenStudy (anonymous):

15/3+21/9 ? How to write in lowest terms

OpenStudy (anonymous):

5/1 +7/3 ?step is next

OpenStudy (dayakar):

multiply first fraction numerator and denominator with 3 and add second fraction

OpenStudy (calculusxy):

\[\large \frac{ 15 }{ 3 } + \frac{ 21 }{ 9 }\] Notice that the denominators are not the same and to add them together we need to make them the SAME. The LCM (least common multiple) will be \(\large 9\) when listing the multiples of \(\large 3\) and \(\large 9\). For the first fraction of 15\3 we need to multiply the numerator and denominator by 3 to get the denominator of 9. And the second fraction of 21/9 the denominator is already 9 so it stays the same. \[\large \frac{ 45 }{ 9 } + \frac{ 21 }{ 9 } = \frac{ 66 }{ 9 } = 7 \frac{ 3 }{ 9 } \text{ }\text{or} \text{ }7\frac{ 1 }{ 3 }\]

OpenStudy (dayakar):

5*3/1*3 +7/3 simplify it

OpenStudy (pawanyadav):

That is 15/3+7/3

OpenStudy (dayakar):

(15+7)/3

OpenStudy (pawanyadav):

I think you know how to add them now also they have same denominator

OpenStudy (calculusxy):

@Pawanyadav I can see where @dayakar is going with this. @dayakar is trying to simply (or put to lowest terms) 15/3 by dividing the numerator and the denominator by 3 to get 5/3 and 21/9 also by 3 to get 7/3. I think that it should result for the same thing? \[\frac{ 5 }{ 1 } + \frac{ 7 }{ 3 } = \frac{ 15 }{ 3 } + \frac{ 7 }{ 3 } = \frac{ 22 }{ 3 } = 7 \frac{ 1 }{ 3 }\]

OpenStudy (calculusxy):

I meant to say 15/3 is going to be 5/1

OpenStudy (pawanyadav):

But with less calculation

OpenStudy (anonymous):

Thanks Q for all of you